The demand for petrol rises from 500 to 600 barrels when the price of a particular scooter is reduced from Rs. 25000 to Rs. 22000. Find out the cross elasticity of demand for the two. What is the nature of their relationship?
1
Expert's answer
2018-09-24T11:25:08-0400
Сross elasticity of demand:
E = (500+600)/(25000+22000)×(600-500)/(22000-25000) = -11/14100
The petrol and the scooter are complementary goods, so cross elasticity of demand is negative